# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ III-B Entropy Analysis of Imitation Learning
Nevertheless, collecting diverse demonstrations is expensive in robotic manipulation [42]. Although huge efforts have been invested in building large-scale manipulation datasets [9, 10], the available quantity of demonstrations is far from sufficient compared with fields like large language models [43]. Currently, the policies trained on these datasets only serve as pre-trained network weights, still requiring task-specific fine-tuning for each downstream application [27, 42]. Thus, exploring strategies to reduce the demonstration data requirement of imitation learning is important.
To conduct this exploration, we first specify that the optimization objective of imitation learning is to approximate the joint observation-action distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$, which describes how to achieve successful manipulation. For notational convenience, we represent the sets of observations and actions in $\mathcal{C}$ as $\mathcal{O}_{\mathcal{C}}$ and $\mathcal{A}_{\mathcal{C}}$. To explore how to reduce the demonstration demand while maintaining policy performance, we need a theoretical framework to quantify the data requirement for approximating $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$. In information theory, the entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ serves as a measure of the amount of information contained within a distribution, reflecting the inherent uncertainty or complexity of data [44]. A higher entropy indicates the distribution contains more information and requires more demonstrations to accurately approximate.
Furthermore, we decompose $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ via the chain rule:
$\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})=\mathcal{H}(\mathcal{O}_{\mathcal{C}})+\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}}).$ (5)
This decomposition reveals that the information entropy of imitation learning consists of two parts, the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. By reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the demonstration volume requirement can be alleviated.
In addition, we expect the manipulation performance not to compromise after reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. To figure out how to realize this goal, it should be first noted that although there are numerous manipulation trajectories meeting constraints, $\pi$ only needs to predict one of them to complete the task. Therefore, when reducing the entropy, the probability of selecting a suitable action leading to successful manipulation should not be decreased. To ensure this, (i) When lessening the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we only remove redundant observation information irrelevant to inferring accurate actions. (ii) During reducing the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the policy should increase the likelihood of predicting a successful trajectory.
Figure 2: This figure illustrates the primary factors affecting the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. The factors include environment, robot, sensor, and task instruction. These factors can be further divided.
III-C Observation Complexity Reduction
To alleviate the demonstration data demand, we first study how to reduce the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, which measures the complexity of the observation set in $\mathcal{C}$. Specifically, we expect to reduce the redundant information in observations that is irrelevant to inferring actions, because removing such information does not hurt action prediction precision. Notably, which kinds of information are irrelevant to actions are not deterministic. They are influenced by how the manipulation system is designed. For instance, if a policy $\pi$ directly predicts the robot hand end-effector movement and then utilizes this movement to control joint rotations based on inverse kinematics, the joint angle observation is unnecessary for inferring actions when the end-effector position observation has been provided. Therefore, the key to reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is devising a manipulation system making most factors irrelevant to inferring actions. Then, these irrelevant factors can be excluded from observations, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
To more thoroughly analyze this problem, we need to know $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is determined by which factors. We illustrate the primary factors in Fig. 2. These main factors contributing to $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ can be categorized into four classes, i.e., environment information, robot proprioception, sensor configuration, and task instruction. Among them, the environment information includes the target object to manipulate, scene background, and ambient conditions (such as lighting brightness and air humidity). These pieces of information can be further divided, like the target object information can be classified into its physical geometry and visual appearance.
Differently, robot proprioception mainly describes the type of robot employed (like Franka Panda or UR5e), where the robot is located in the environment, and the action space (like end-effector control or joint control). Sensor configuration specifies what kinds of sensors are adopted (like RGB camera and depth sensor), how the sensors are deployed in the environment (the number of sensors and their poses), and the parameters of these sensors. A task instruction specifies the task target. It can be represented in different modalities, including text [11], image [40], voice [45], etc.
Summarily, the above four classes of factors play distinct roles in shaping the complexity of the observation set $\mathcal{O}_{\mathcal{C}}$. To reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we need to design a system making most factors irrelevant to predicting actions. We explain how such a specific system is devised in Section IV.
Figure 3: The sub-figure (a)-(c) illustrate the procedures that a robot hand presses a button. We can find that if the policy input does not include historical memory and moving velocity, the policy observations are the same between sub-figure (a) and (c), in which the robot hand is approaching the button and moving away from the button, respectively.
III-D Action Prediction Ambiguity Reduction
In this subsection, we discuss how to reduce $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ in Eq. (5), which represents the action prediction ambiguity given observations. This ambiguity is mainly caused by the multimodal nature of robotic manipulation, meaning there exist multiple valid action trajectories given the same observation [46]. For example, in the task of grasping a block, a robot arm can complete the task by approaching the block from either the left or the right side. Moreover, the ambiguity becomes more pronounced when the observation information is insufficient. For instance, Fig. 3 illustrates an example that a robot hand is pressing a button. In Fig. 3 (a)-(c), the robot hand is approaching the button, pressing the button, and departing from the button, respectively. If the policy input only includes the current visual observation and robot joint angles, the observations are the same between Fig. 3 (a) and (c), while the accurate end-effector moving directions are different.
Hence, larger action prediction ambiguity not only increases demonstration demand, but also raises the risk of making incorrect decisions. To handle this problem, more comprehensive information supporting accurate action prediction should be provided. According to the example in Fig. 3, incorporating historical memory into policy observation is effective, especially for complex and long-chained manipulation tasks. This is because historical memory enables the policy to know which manipulation steps have been completed.
However, previous works often find incorporating historical memory degrades success rates [8]. One key reason is that these works choose to model memory using historical action trajectories. As analyzed in VIP [40], when the attribute to predict (the future action trajectory in this case) and a part of the policy input (the historical action trajectory) fall in the same modality, the policy often chooses to overlook the input information in other modalities (like images). This is because it is easier to predict future actions based on historical actions than information in different modalities. A learnable network tends to choose the easiest way of regressing future actions while overlooking the exploitation of information in other modalities. This problem is termed as shortcut learning and has been well analyzed in previous literature [47].
An effective strategy for handling the shortcut learning problem is preventing using input in the same modality as output. In this case, it means avoiding modeling memory with historical action trajectories. Drawing on this insight, we represent memory with images. However, directly incorporating historical images into policy input leads to a huge computational burden [48]. To address this problem, we develop an efficient recurrent encoder for modeling visual memory [49], which is explained in Section IV-E. By utilizing historical memory, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is reduced significantly.
Figure 4: As shown in the sub-figure (a), in the task “pick up a yellow square”, the accurate actions present a multimodal distribution due to the existence of two yellow squares. In the sub-figure (b), if the policy is an unimodal distribution estimator, the action corresponding to the minimum expected error is towards the pink circle, an incorrect moving direction.
III-E Multimodal Distribution Estimation
As described in Section III-B, imitation learning optimizes a policy $\pi$ to capture the probability distribution inherent in demonstration data. Due to the action prediction ambiguity, the expert demonstration distribution is multimodal. Therefore, $\pi$ should be capable of fitting a multimodal distribution.
Recently developed policies mostly regress future actions directly, and the policy network parameters are updated based on loss functions like mean absolute error [8]. Such loss function designs imply these policies regress to unimodal distributions with fixed uncertainties [50]. In other words, these policies are unimodal distribution estimators. In Fig. 4, we show what happens when we employ an unimodal distribution estimator to fit a multimodal distribution. Specifically, the robot needs to pick up a yellow square, and the valid actions present a multimodal distribution due to the existence of two yellow squares. The optimization objective is to minimize the regression error between the multimodal demonstration distribution and unimodal policy prediction, which is formulated in Eq. (4). Similar to the process of mean regression [51], this optimization process leads to an action prediction taking the mean of all valid actions (the mean expected value of this multimodal distribution). As visualized in Fig. 4 (b), the robot moves towards an incorrect target, the middle position between these two yellow squares.
To handle this problem, we propose to model the policy output as Gaussian Mixture Distribution [52] with $K$ components, which can naturally represent multimodal distributions. What the policy predicts is the means, variances, and weights of these Gaussian components, denoted as $\mu_{k}$, $\sigma^{2}_{k}$, and $\alpha_{k}$ for the $k_{\rm th}$ component, respectively. Thus,
$\pi(a_{t}|o_{t},\tau)=\sum\limits_{k=1}^{K}\alpha_{k}\mathcal{N}(\mu_{k},\sigma^{2}_{k}).$ (6)
The component weights $\{\alpha_{k}\}_{k=1}^{K}$ satisfy $\sum\limits_{k=1}^{K}\alpha_{k}=1$ and $\alpha_{k}\geq 0$ for all $k\in\{1,2,\cdots,K\}$. We update the policy following the optimization objective in Eq. (3). In this way, our developed policy is able to fit complex demonstration data with multimodal distributions.
Summarily, the content in this subsection ensures that we can approximate the multimodal demonstration distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ accurately, and the analysis from Section III-B to Section III-D enables a policy to learn $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ with fewer demonstrations. Guided by these insights, we develop a generalizable manipulation system GEM, which is elaborated in Section IV.
IV Generalizable Entropy-based Manipulation
In this section, we introduce our developed dynamic object manipulation system GEM. GEM is devised following the theoretical framework in Section III. Only using hundreds of manual demonstrations and thousands of programmatic demonstrations collected from simulated environments, GEM can generalize to diverse unseen target objects, moving dynamics, background environments, and robot embodiments.
Figure 5: Our built hardware platform and remote teleoperation handle. Specifically, Sub-figure (a) illustrates the real robot environment for studying dynamic object manipulation. GEM controls the robot arm to perform diverse tasks given the observations from the external and wrist RGB-D cameras. Sub-figure (b) shows our designed remote handle for teleoperating the robot arm to collect demonstration data.
IV-A Hardware Setup
To study the dynamic object manipulation problem, we build a hardware platform as shown in Fig. 5. In this platform, objects are driven by a conveyor belt at an adjustable velocity. A robot arm with a gripper is mounted on a side of this conveyor belt to manipulate the target object. Two RGB-D cameras are deployed to capture environment observations, one on the robot arm (the wrist camera) and the other on a fixed third-person perspective bracket (the external camera). GEM controls the robot arm to manipulate objects based on the observations from these two cameras.
As illustrated in Fig. 5 (b), a teleoperation handle is designed to teleoperate the robot arm in Fig. 5 (a). Specifically, we control the motion of the robot arm end-effector based on the relative pose change of the handle derived from the SLAM camera. The opening and closing of the gripper is regulated through the trigger pull degree. A battery is mounted on the handle to power all the devices in this handle. The control signals from the SLAM camera and trigger are wirelessly transmitted to the robot arm by the main control board, thereby enabling flexible remote teleoperation. The teleoperation handle is only utilized to collect demonstration data for training the learnable network, which we explain in Section IV-E. The handle is not used when testing the method.
Figure 6: The overall pipeline of GEM. GEM primarily consists of two blocks, the object-centric geometry observation and hybrid action control. In object-centric geometry observation, we exclude the many factors irrelevant to predicting actions from policy observation, and thus the observation entropy is reduced. In hybrid action control, we combine classic visual servo control with learnable network control. In the learnable network control, we design the recurrent probabilistic network, which models historical vision observations as memory. In this way, the action prediction entropy is decreased. Only using hundreds of demonstrations for training, GEM can be generalized to diverse unseen scenes. In this figure, the pink modules are implemented based on learnable neural networks and the yellow modules are non-learnable algorithms.
IV-B GEM Overall Framework
The overall framework of GEM is illustrated in Fig. 6. It can be observed that GEM comprises two blocks, i.e., object-centric geometry perception and hybrid action control. The object-centric geometry perception block extracts 3D point clouds of the target objects in the external and wrist sensor views. The design of this block follows the principles summarized in Section III-C, which points out the importance of excluding factors irrelevant to predicting actions from observations. In this way, the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is decreased and less data is required to achieve strong generalization.
In hybrid action control, we combine a classic visual servo control algorithm [1] and an imitation learning based neural network control algorithm to compute the robot end-effector trajectories. Specifically, the visual servo control algorithm is responsible for tracking the target object to manipulate, and the neural network control algorithm (the recurrent probabilistic network) generates manipulation actions of different skills according to input skill type instructions. The robot end-effector action is the sum of the actions produced by these two algorithms. This design allows the recurrent probabilistic network not need to consider the moving dynamics (like velocity and acceleration) of the target object, further reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. In addition, the developed recurrent probabilistic network can model long-term historical memory, thereby decreasing the action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. The output actions of the recurrent probabilistic network are represented in Gaussian Mixture Distributions as described in Section III-E. Hence, the manipulation multimodal ambiguity problem is addressed.
To enable GEM to adapt to different robot hand embodiments, the hybrid action control block predicts the translation and orientation of the robot end-effector. Based on this prediction, we calculate the corresponding robot joint rotations via inverse kinematics [53], and these rotations are utilized to operate the movements of robot arms. This strategy allows for compatibility with various robot embodiments.
IV-C Object-Centric Geometry Perception
This subsection describes a perception module that produces observations with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. To design this module, we first identify the redundant factors in observations. If these factors do not provide necessary information for predicting actions, we can exclude them from observations to reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ while maintaining manipulation performance. Fig. 2 summarizes the primary factors affecting $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. By analyzing them, we first observe that successful manipulation mainly depends on the geometric structures of objects rather than their appearances. However, most embodied networks take images as input [46], which conveys visual appearance directly but lacks explicit 3D geometry. To handle the missing geometry, we employ uncolored 3D points as policy input. These 3D points can be easily accessed via the RGB-D cameras shown in Fig. 5. Besides, some recent deep learning models can produce promising dense depth maps from monocular or binocular images [54, 55], which facilitates obtaining depth. In addition, 3D points provide representation less affected by sensor parameters. This is because 3D points exist in the real world and are not represented in the pixel coordinate system.
Since environment background does not affect manipulation interaction, we adopt a lightweight instance segmentation model [56] to segment points of concerned objects and remove background points. Such an instance segmentation model is readily obtainable due to the availability of many open-world segmentation models [57]. Notably, multiple objects could be observed in a single frame. For example, in the “pick up the spoon out of the cup” task illustrated in Fig. 6, the 3D points of both the spoon and cup are important. In industrial applications, as manipulation procedures are predetermined, we can define which objects to segment in advance. For tasks with unfixed procedures, large language models [26] could be employed to reason about which objects to focus on.
Notably, a manipulation policy needs to classify the points belonging to different objects, like the cup and spoon in the example shown in Fig. 6. Nevertheless, using uncolored 3D points as policy input makes this classification challenging. To address this obstacle, we leverage instance segmentation results to append distinct instance identifiers (IDs) to the points of various objects. With the instance IDs, the object classification problem is addressed. Summarily, combining the above designs, we propose to use uncolored 3D points of concerned objects appended with corresponding instance IDs as the observation representation with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
IV-D Hybrid Action Control
This subsection explains how to further reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ by devising the hybrid action control block in Fig. 6. The basic idea is to leverage the complementary strengths of classic control and learnable network control. Classic control algorithms, such as visual servo control [1], excel at stably and rapidly guiding the robot arm to approach and track moving objects. However, they often struggle with collision-rich interactions, such as grasping, inserting, and rotating objects. This is because of the complexity of modeling contact dynamics. On the other hand, learnable network control methods, particularly those grounded in imitation learning, are adept at handling such interactions by capturing the underlying patterns in expert demonstrations. By combining these two approaches, we enable the learnable network to concentrate on generating collision-rich interactions, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
Figure 7: This figure illustrates how classic control and learnable network control are combined. Specifically, classic control is utilized to approach and track the moving target object. The learnable network control starts generating manipulation actions once stable tracking is achieved. The grippers and target objects marked in gray, dark yellow, and reddish brown correspond to their positions at Moments 1, 2, and 3, respectively. The pink gripper indicates the position that would be achieved using classic control alone at Moment 3.
As described in Section IV-C, the object-centric geometry perception module extracts the 3D points of the target object to manipulate from both the external and wrist sensors. Denote the two groups of 3D points as $\{(x_{i}^{e},y_{i}^{e},z_{i}^{e})\}_{i=1}^{N^{e}}$ and $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, respectively. The classic control, consisting of an object state estimator and an end-effector tracking planner as shown in Fig. 6, operates the robot hand to track the moving target object based on the external view 3D points. Specifically, the object state estimator first computes the top centroid of the external view 3D points, denoted as $(x_{c}^{e},y_{c}^{e},z_{c}^{e})$, where $x_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}x_{i}^{e}$, $y_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}y_{i}^{e}$, and $z_{c}^{e}=\max\limits_{i=1}^{N_{e}}z_{i}^{e}$ (the positive z-axis points upward). Based on the top centroids at different timestamps, we can estimate the target object velocity $(v_{x},v_{y},v_{z})$ using Gaussian process regression [58] implemented in the object state estimator. If the target object is occluded in the external view camera, its real-time top centroid is also predicted by this estimator based on historical centroid records using Gaussian process regression.

### v2

@@ III-C Imitation Learning
III-D Optimization Path Shaping
The optimization objective of imitation learning is solely to minimize the cost value calculated by the difference measurement function defined in Eq. (4) and does not specify which observation feature to exploit. As discussed in Section III-B, the observation feature in colored point clouds can be categorized into the appearance and geometry information. If an RGB-D policy learns to only use appearance information, its learning process becomes similar to an image-only policy, like ACT [60]. In this situation, the policy needs to infer the 3D geometry of objects from visual appearance implicitly to estimate accurate manipulation poses [54]. By contrast, if a policy primarily relies on geometry, it can predict manipulation poses in a more direct and precise manner.
As discussed in Section III-B, the appearance gap between simulated and real environments motivates this work to shape the policy optimization path to prioritize the geometric information, which is more consistent between simulated and real-world observations. To analyze how to achieve this target, we draw on the deep learning theory of feature predictivity established in prior literature [17]. In this theory, predictivity basically means how useful a feature is for accurately inferring training labels, and a policy tends to pick features with higher predictivity to minimize training loss.
Since the geometry of an object directly reflects its 3D structure, geometry is inherently of higher predictivity for estimating manipulation actions. However, in simulation, scene layout is often not diverse, e.g., objects with specific textures always appear in fixed areas. This makes appearance features (e.g., color, texture) accidentally gain unexpected high predictivity. Correspondingly, the policy overfits to these scene-specific appearance patterns, using them to predict actions instead of the generalizable geometry.
To prioritize geometry information and alleviate appearance overfitting, we propose to reduce the predictivity of appearance features by introducing visual noise, thereby shaping the policy optimization path towards geometry. However, persistent appearance noise would prevent the policy from leveraging useful appearance cues. For instance, in block insertion tasks where a block should be placed into a container, appearance helps distinguish these two objects, and relying solely on geometry could lead to misidentification.
To resolve this appearance utilization trade-off, we propose appearance noise annealing, a strategy that gradually removes appearance noise during training. Specifically, strong appearance noise is applied at the beginning of training to guide the policy to focus on geometry, followed by steady noise reduction until the noise is fully removed. Since the policy already learns to use geometry for loss minimization in early training stages, it will keep prioritizing geometry after noise is removed. For scenarios where geometry alone is insufficient (e.g., distinguishing closely placed distinct objects), the policy can learn to exploit useful appearance texture after the appearance feature becomes obvious.
IV Policy Design
This section presents the proposed manipulation policy, GEM. Specifically, Section IV-A introduces the overall framework of GEM. Then, Section IV-B and Section IV-C explain the developed appearance noise annealing and manipulation action decomposition strategies separately. Afterwards, Section IV-D describes the network implementation details in GEM. Finally, Section IV-E elaborates how we build simulation environments of dynamic object manipulation on conveyor systems to generate demonstration data efficiently.
Figure 3: The overall framework of GEM. The policy observation is the colored 3D points captured by RGB-D cameras. The appearance noise annealing strategy is utilized to perturb the color feature in policy observation. Based on this observation, the GEM network predicts a multi-modal action distribution represented as a Gaussian Mixture Model (GMM), a point segmentation mask of the target object to manipulate, and a status flag marking the completion of manipulating one object. We generate interaction actions and tracking actions from the network outputs and combine them to control the robot end-effector.
IV-A Policy Framework
We illustrate the overall framework of the GEM policy in Fig. 3. To ensure sufficient information in the policy environment observation, we mount two RGB-D cameras in the dynamic object manipulation system. Specifically, an external-view camera is installed opposite the robotic arm, with its perspective directed at the conveyor belt. Another wrist-view camera is affixed to the wrist of the robotic arm. The colored 3D points perceived by these two cameras are merged as a single point cloud to provide environment observation.
As mentioned in Section III-D, we design an appearance noise annealing strategy that shapes the policy optimization path by applying strong color augmentation to the point cloud at the beginning of training and gradually removing the noise until the training end. This strategy guides the policy to focus more on the 3D geometry of objects. Next, as shown in Fig. 3, the augmented point cloud is input to the GEM network, which consists of Transformer encoders, decoders, and a memory module. The memory module represents the historical observation information as an implicit embedding and updates it in a recurrent manner. Incorporating memory information into the network input helps the network be aware of the completion progress of manipulation.
The prediction produced by the GEM network includes an action probabilistic distribution, a point segmentation mask, and a status flag as visualized in Fig. 3. The action probabilistic distribution is modeled using a Gaussian Mixture Model (GMM) [41]. Unlike the popular implementations of recent policies that directly estimate deterministic actions [13], the GEM network predicts a probabilistic distribution over actions, as such an output naturally captures the inherent indeterminism of manipulation decisions. The experimental results suggest that this design can effectively alleviate the multi-modal ambiguity in imitation learning and is more efficient than the commonly employed strategies like diffusion process [7] and flow matching [28]. During inference, we decode the interaction actions from the GMM distribution by selecting the action with the maximum probability value at each timestamp.
The point segmentation prediction is to segment out the 3D points belonging to the target object. This prediction has two benefits, i.e., guiding the policy to focus on the target object and generating tracking actions. For the first benefit, we expect the GEM network to concentrate on the 3D geometry of the target object rather than the whole environment, as the environment background has little effect on successful manipulation. We can understand this guiding process from the view of optimization path shaping, since it increases the importance of target object points on minimizing training loss. The second benefit is we can approximate the 3D centers of target objects based on the segmentation results and make the robot end-effector track the target objects moving in various velocities. We realize this dynamic object tracking by decoding the point cloud segmentation results into tracking actions.
Both the interaction actions and tracking actions represent the relative pose offsets of the robot end-effector. By combining them following our developed algorithm based on the predicted status flag, we obtain the manipulation actions utilized to control the rotations of robot joints based on inverse kinematics [25]. This design of predicting end-effector relative offsets enables the policy to generalize across different robot embodiments. Refer to Section IV-C for the details of how we generate and utilize the predicted actions.
IV-B Appearance Noise Annealing
Figure 4: In the appearance noise annealing strategy, strong color perturbation (noise ratio is 1.0) is applied to the input point cloud at the initial phase of training. As the training progresses, the perturbation is gradually removed.
As discussed in Section III-D, we develop a simple yet effective strategy to shape the policy optimization path, enabling it to focus on object geometry and generalize to real-world tasks via simulated demonstrations. In this strategy, we apply strong color perturbation to the input point cloud at the initial stage of training. According to the aforementioned theory of feature predictivity, this operation reduces the feature predictivity of 2D appearance, thereby compelling the GEM network to prioritize 3D geometry in minimizing training loss. Then, with the progress of training, the color perturbation is gradually weakened. Since the network has already learned to predict manipulation actions using the observed 3D geometric structure of the environment, it does not revert to over-relying on visual appearance after the color perturbation is removed. Furthermore, after the visual features become prominent, the network begins to learn how to leverage visual appearance to handle tasks that cannot be adequately addressed by 3D geometry alone, such as distinguishing two adjacent objects. In the following, we elaborate on the implementation details of how the color perturbation is applied.
Denote the input point cloud at timestamp $t$ with $N$ points as $Q^{t}=\{q_{i}^{t}\}_{i=1}^{N}$, where $q_{i}^{t}=(x_{i}^{t},y_{i}^{t},z_{i}^{t},r_{i}^{t},g_{i}^{t},b_{i}^{t})$. In $q_{i}^{t}$, $(x_{i}^{t},y_{i}^{t},z_{i}^{t})$ is the point coordinate in the wrist-view camera coordinate system, and $(r_{i}^{t},g_{i}^{t},b_{i}^{t})$ represents the visual color of this point. The color channels have been normalized to the interval $[0,1]$, i.e., $(r_{i}^{t},g_{i}^{t},b_{i}^{t})\in[0,1]^{3}$. To disturb the color of the point $q_{i}^{t}$, we sample three independent random scalars $\eta_{i,r}^{t},\eta_{i,g}^{t},\eta_{i,b}^{t}\sim\rm{Uniform}(0,1)$ and assemble them into an RGB noise vector denoted as $\eta_{i}^{t}=(\eta_{i,r}^{t},\eta_{i,g}^{t},\eta_{i,b}^{t})\in[0,1]^{3}$.
To control the strength of the noise perturbation $\eta_{i}^{t}$ applied to $(r_{i}^{t},g_{i}^{t},b_{i}^{t})$, we define a noise ratio factor $\gamma_{\xi}$ that varies with training progress, where $\xi\in[0,1]$ denotes the normalized training progress. To ensure a strong perturbation at the beginning of training and make the perturbation gradually vanish as learning proceeds, the variable $\gamma_{\xi}$ is defined as:
$\gamma_{\xi}=\begin{cases}1,&\xi\in[0.0,0.2),\\ 1-\sigma(k(\xi-m)),&\xi\in[0.2,0.8],\\ 0,&\xi\in(0.8,1.0],\end{cases}$ (5)
where $\sigma(\cdot)$ is fomulated as $\sigma(x)=1/(1+e^{-x})$, and $k$ and $m$ are set to 10 and 0.5, respectively. The curve of $\gamma_{\xi}$ evolving with the training progress is illustrated in Fig. 4. The point color $c_{i}^{t}=(r_{i}^{t},g_{i}^{t},b_{i}^{t})$ is augmented into the perturbed vector $\tilde{c}_{i}^{t}$ following:
$\tilde{c}_{i}^{t}=\gamma_{\xi}\eta_{i}^{t}+(1-\gamma_{\xi})c_{i}^{t}.$ (6)
We can observe from Fig. 4 that the appearance observation of the environment gradually becomes obvious with the progress of training, which satisfies the algorithm design requirement of optimization path shaping. The point cloud with the perturbed color attributes $\{\tilde{c}_{i}^{t}\}_{i=1}^{N}$ serves as the direct input to the GEM network. If $\gamma_{\xi}$ is 1, the visual feature observed from the point cloud is pure random noise. When $\gamma_{\xi}$ becomes 0, the appearance observation is noise-free.
IV-C Manipulation Acition Decomposition
As illustrated in Fig. 3, GEM first decodes the network output into the tracking and interaction actions and then combines them as the manipulation actions. Thanks to this strategy, GEM can generalize to manipulating objects moving in diverse unseen velocities, even though the GEM network is trained solely on static object demonstrations. This capability makes GEM practical for real-world industrial manipulation on conveyor belts, where the belt speed varies. To explain this control strategy clearly, we first describe how to derive the tracking actions in the following.
Figure 5: This figure illustrates how the predicted tracking actions and interaction actions are integrated to perform dynamic object manipulation. Specifically, tracking actions are responsible for approaching and tracking the moving target object. The interaction actions begin to take effect once stable tracking is achieved, and they are for conducting collision-rich contact with the target object, such as picking and rotating.
Based on the point segmentation mask from the GEM network, we extract the spatial coordinates of the 3D points belonging to the target object to manipulate, denoted as $\{(x_{i}^{\tau},y_{i}^{\tau},z_{i}^{\tau})\}_{i=1}^{N^{\tau}}$, where $N^{\tau}$ is the total number of 3D points. The top centroid of these points is defined as $(x^{c},y^{c},z^{c})$, where $x^{c}=\frac{1}{N^{\tau}}\sum\limits_{i=1}^{N^{\tau}}x_{i}^{\tau}$, $y_{c}=\frac{1}{N^{\tau}}\sum\limits_{i=1}^{N^{\tau}}y_{i}^{\tau}$, and $z_{c}=\max\limits_{i=1}^{N^{\tau}}z_{i}^{\tau}$ (the positive z-axis points upward). By tracking this top centroid across time, we can derive the translational velocity $(v^{\tau}_{x},v^{\tau}_{y},v^{\tau}_{z})$ of the target object with simple estimation algorithms like Gaussian process regression [50].